OpenStudy (anonymous):

6/7 * 9 in a fraction form

OpenStudy (anonymous):

54/7

OpenStudy (anonymous):

Remeber BIDMAS so strictly division comes first then multiplication which gives \[\frac{6}{7}*9=7.714 \] . but this can also be written with the same result as \[\frac{6*9}{7}=7.714\]and as above \[\frac{54}{7}=7.714\]